Force between two identical sphere charged for the same charge F. If 50%charge of one sphere is transferred to another sphere then now the new force will be

Grade:12

4 Answers

Arun
6 years ago
Dear student
 
In the starting
F = K Q²/d²
Now
F' = k (3Q/2) (Q/2) / d² = 3kQ²/4 d²
Hence
F' = 3F/4
Regards
Arun (askIITians forum expert)
Krish Gupta
3 years ago
Let’s say the initial charge be Q on both the charges and let distance be d, then force F
F = kQ²/d²
Now, when half the charge has transffered 
F' = k (3Q/2) (Q/2) / d² = 3kQ²/4 d²
Hence, new force is
F' = 3F/4
 
Thanks, hope it helps.
ankit singh
3 years ago
 
F = kQ²/d²
Now, when half the charge has transffered 
F' = k (3Q/2) (Q/2) / d² = 3kQ²/4 d²
Hence, new force is
F' = 3F/4  ..thank u regards ankit singh
Rajeev Siyag
2 years ago
F = kQ²/d²
Now, when half the charge has transferred
F' =k(3Q/2)(Q/2)/d² = 3kQ²/4d²
Hence, new force is
F' = 3F/4
You can use Q = q ) 
Thanks